Can you also show how you calculated the answers? Mostly stuck on the answer to NWC and how to estimate cash flows (and everything after that)Conch Republic Electronics Conch Republic Electronics is a midsized electronics manufacturer located in Key West, Florida. Th
Answer #1

1). Payback period = 2.698 years

2). Profitability index = 1.515

3). IRR = 28.07%

4). NPV = 19,845,464.30

All calculations show in the table below:

2 Formula Equipment cost 38500000 TR = u*p VC=u*vc Year (n) Initial investment (11) Unit sales Price/unit Variable cost/unit

Free Cash Flow calculation: NWC = 20%*TR Net Working Capital 659200010918000 89610008034000 5562000 NWCn-1 - NWCn; Recovery o

5).

5). Sensitivity of NPV to change in price of the smartphone can be found by varying the price to get the new NPV.

NPV = 19,845,464 when price per unit = $515

NPV = 18,765,838 when price per unit = $510

Sensitivity of NPV to price = (new NPV - old NPV)/(new price - old price) = (18,765,838 - 19,845,464)/(510-515) = 215,925.32

6). Sensitivity of NPV to the quantity sold:

NPV = 19,845,464 when original quantities are sold

NPV = 20,669,521 when original quantities per year are increased by 1,000

Sensitivity of NPV to quantity = (new NPV - old NPV)/change in number of units sold per year = (20,669,521-19,845,464)/1,000 = 824.06

7). Since NPV is positive, the project is profit making, so Conch Republic should produce the new smartphone.

8). If other models register lower sales due to introduction of the new model then the loss is sales will be part of the NPV analysis.